OpenStudy (anonymous):

can anyone tell me what I am doing wrong? I have the formula and answer but when I try to plug in the formula in my calculator its not the same!!

OpenStudy (anonymous):

R = P i / (1 - (1+i)^(-n)) ... R = payment , P = loan principal, i = interest rate, n = number of periods 1100 = P (0.075/12) / (1 - (1 + 0.075/12)^(-30*12)) P = 157319.39 1400 = P (0.075/12) / (1 - (1 + 0.075/12)^(-30*12)) P = 200224.68 Answer: $157,319.39 to $200,224.68

OpenStudy (anonymous):

what's wrong?

OpenStudy (anonymous):

are u ok with this one?